OpenStudy (monica0597):

(x-2)^1/3=5 can anyone explain how to do this?

OpenStudy (solomonzelman):

\[\huge\color{blue}{ (x-2)^{(1/3)}=5 } \] we know: \[we~~~~know:~~~~~~~~~\huge\color{blue}{ A^{(b/c)}=\sqrt[c]{A^b} } \] \[So,~~~~~~~~~\huge\color{blue}{\sqrt[3]{(x-2)^1}=5 } \]\[~~~~~~~~~~~~~\huge\color{blue}{\sqrt[3]{(x-2)}=5 } \] PAY CLOSE ATTENTION: \[~~~~~~~~~~~~~\huge\color{blue}{\sqrt[3]{(x-2)}=\sqrt[3]{5^3} } \] \[~~~~~~~~~~~~~\huge\color{blue}{x-2=5^3 } \] do it from here.

OpenStudy (monica0597):

all i see is this: we know: \[we~~~~know:~~~~~~~~~\huge\color{blue}{ A^{(b/c)}=\sqrt[c]{A^b} } \] \[So,~~~~~~~~~\huge\color{blue}{\sqrt[3]{(x-2)^1}=5 } \]\[~~~~~~~~~~~~~\huge\color{blue}{\sqrt[3]{(x-2)}=5 } \] PAY CLOSE ATTENTION: \[~~~~~~~~~~~~~\huge\color{blue}{\sqrt[3]{(x-2)}=\sqrt[3]{5^3} } \] \[~~~~~~~~~~~~~\huge\color{blue}{x-2=5^3 } \]

OpenStudy (solomonzelman):

Do you see? \[~~~~~~~~~~~~~\huge\color{blue}{(x-2)^{(1/3)}=5 } \]

OpenStudy (monica0597):

yep

OpenStudy (solomonzelman):

Ok, do you understand this process in blue?

OpenStudy (monica0597):

i see everything in black but i got the answer. i saw the work you showed me. thank you :)

OpenStudy (anonymous):

laTex is not working on my computer DX

OpenStudy (solomonzelman):

You welcome!

OpenStudy (solomonzelman):

@souring, what browser are you using?

OpenStudy (solomonzelman):

sourwing

OpenStudy (anonymous):

chrome

OpenStudy (anonymous):

it worked before

OpenStudy (solomonzelman):

yeah, I started recently seeing people that have problem with latex.

OpenStudy (solomonzelman):

it could depend on various things.